# The physical basis of teleological evolutions

### v6

@@ 3.1 Updating the EPR critique of the quantum description
Of course, the first assertion that the quantum description is incomplete goes back to the famous EPR paper by Einstein, Podolsky and Rosen $\left[28\right]$. As is well known, Einstein in particular, argued that the ordinary quantum description of quantum nonlocality is incomplete because it does not physically explain the “spooky action at a distance” that it mathematically predicts. They also stressed the need to devise ”hidden variables” that would complete the quantum description.
@@ 3.1.2 Generalization
In the first place, to be a quantum superposition of causal loop, the quantum algorithm does no more need to be optimal444 If the solution is not an invertible function of the problem setting, we can have one-to-one correlation between the initial and final measurement outcomes by requiring that the final measurement measures both the solution and the problem-setting. . The time-symmetrization of its ordinary quantum description always yields that it is as if the problem-solver knew in advance one of the possible halves of the information about the solution. In equivalent terms, there is a corresponding reduction of the computational complexity of the problem to be solved, and the dimension of the Hilbert space in which the search for the solution (or, identically, the one-to-one correlation) takes place. However, any of these facts just creates the room for a quantum speedup. To also have a quantum speedup, it takes that the problem-solver, or more generally the unitary evolution, exploits such a room. The room is fully exploited in the optimal quantum case, but it can also be not exploited at all or only partly exploited.
@@
By the way, as is well known, the EPR paper inspired works of great fundamental and practical interest like Bell’s inequality $\left[33\right]$ and Ekert’s quantum cryptography based on Bell’s theorem $\left[34\right]$. We do not understand why the argument concerning the incompleteness of the quantum description, central in that paper, has been removed by the scientific community. As is also well known, the question of the incompleteness of the quantum description has been at the center of the famous Einstein-Bohr dispute. Bohr, who was for the completeness, was eventually decreed the winner by the scientific community. Apparently that was once and for all.
@@ 3.2 Formalization
(2) This point is propaedeutic to point (4). It shows how to relativize with respect to the problem-solver the ordinary quantum description of the quantum algorithm, which is with respect to the customary observer external to the quantum process. Of course, the problem-solver must be concealed from the problem-setting selected by the initial measurement, which would give her full information about the solution of the problem555 We must assume that the problem-solver knows everything about the structure of the problem, thus also the relation between problem-setting and solution. .
@@ 3.2.1 Extending the usual quantum description of quantum algorithms to the process of setting the problem
We should start from the quantum state in the upper-left corner of Table I diagram and then follow the vertical and horizontal arrows. Initially, at time $t_{1}$, both registers, each by itself, are in a quantum superposition of all their basis vectors. By measuring $\hat{B}$ (the content of $B$) in this initial state, Bob selects the number of the drawer with the ball at random, say it comes out $01$ – see the state of register $B$ under the vertical arrow666 By the way, Bob would be free to unitarily change at will the sorted out number of the drawer with the ball. We omit this operation that would change nothing. .
@@
Note that this relativization relies on relational quantum mechanics $\left[35,36\right]$, where the quantum state is relative to the observer. In the following, we will continue to call the quantum description with respect to the external observer the ordinary quantum description of the quantum algorithm. Instead we will always specify that the other quantum description is that with respect to the problem-solver.
@@ 4 The possible physical basis of natural teleological evolutions
In the universe we are in, the search for the value of life becomes search for life, of course endowed with the same quantum speedup. We have also seen that, as a consequence, the evolution of life on Earth, obtainable from that of the universe by tracing out all the rest, even if decohered, would inherit the quantum speed up of the evolution of the universe toward life and would therefore be teleological too. The same would necessarily hold for the evolutions of species, etc..
@@ 4.2 Decohered teleological evolutions: what would they look like?
The above explanation is puzzling, of course. We should keep in our mind that we are under the quantum cosmological assumption that should make macroscopic evolutions, the partial trace over all the rest of a quantum evolution of the universe, as puzzling as the microscopic quantum world.
@@ 4.3 How to model teleological natural evolutions
Let us note that, since the liquid water stage of it, the evolution of life on Earth should have been almost completely independent from the evolution of the rest of the universe. If it were completely independent, it would be factorizable out of the evolution of the universe, thus itself quantum and endowed with quantum speedup. Therefore, an obvious, approximate, way of modeling the decohered evolution of the living on Earth endowed with quantum speedup is just to ignore the fact it is decohered and model it in a quantum way. There are encouraging precedents to this. For example, it has been shown that modeling in a quantum way the interplay between species of predators competing for the same preys allows to explain the empirically ascertained species’ asymptotic behavior in a way otherwise difficult to explain $\left[37\right]$.
@@ 4.4 Impact on the natural sciences
This tension is captured by the following sentence by the evolutionary biologist Haldane $\left[38\right]$:
@@
No biologist, of course, does not mean no scientist. A paradigmatic example is the American philosopher Nagel $\left[39\right]$, who argues that the principles that explain the emergence of life should be teleological, rather than materialistic or mechanistic in character.
@@
Summing up, given the above, the impact of the present teleological view of natural evolutions on today’s evolutionary biology would seem to be almost a head on collision.
Almost because, on the one side, in recent years a group of evolutionary biologists $\left[40\right]$ expressed the need to rethink Darwin’s theory of evolution. That was after the discovery that some species’ evolution steps occurred far too quickly to be solely ascribed to random changes in the species’ DNA pool. Of course, the present teleological view of evolutions, with the associated quantum speedup, would offer a way out.
On the other side, there are scientists who have began to apply essentially quantum mechanical features like quantum entanglement and nonlocality to the macroscopic game of species, showing that this can explain some empirically ascertained features of it otherwise difficult to explain $\left[37\right]$.
@@ 4.5 Impact on philosophy
Of course, the current theory of teleological natural evolutions has a precise conceptual precedent in a very ancient notion of natural philosophy $\left[41\right]$. We should therefore be allowed to look for other possible relations between this theory and philosophy. As we are going to see, their abundance is a surprising interdisciplinary aspect of the present work.
@@
Let us call the present teleological view of evolutions and the causal view of today’s physics and the natural sciences respectively the mutual causality view777 Mutual causality between the initial and final measurement outcomes is what generates the quantum superposition of the time-symmetric causal loops. and the causal view.
The causal view and the mutual causality view seem to have respectively subtended two opposing views throughout the history of philosophy. The causal view would underlie the materialist, mechanistic and reductionist views, while the mutual causality view would underlie the idealist, teleological and monist views respectively.
Consider the opposition between materialism and idealism. The former, about the observer being causally created by physical reality, is a causal view; the latter, about the mutual creation between observer and physical reality, like in the participatory Anthropic Principle of Wheeler $\left[31\right]$, and the anthropic cosmological principles of Barrow and Tipler $\left[12\right]$ is a mutual causality view.
@@
(i) For Plato, “Ideas [of our mind] are objective perfections that exist in themselves and for themselves and, at the same time, they are the causes of natural phenomena and constitute their unity” $\left[42\right]$. In modern language, the cause of natural phenomena that constitute their unity are the fundamental physical laws. The identity between the fundamental physical laws since the beginning of time and the ideas of our mind today is formally reminiscent of the identity (and mutual causality) between the number of the drawer with the ball (the problem-setting) randomly selected by Bob’s measurement at the beginning of the quantum algorithm and the number of the drawer with the ball (the solution of the problem) selected by Alice’s measurement at the end of it. It is as if the number of the drawer with the ball was determined by both the problem-setter and the problem-solver.
@@
The above should also go along with Fritjov Capra’s $\left[43\right]$ notion that there is identity between the fundamental states of consciousness described by eastern theosophies and the fundamental laws of quantum physics.
(ii) We do not need to be anthropocentric. This mutual causality would also be between physical laws and spatial intelligence, often superior to our in non-human animals. The assumption that spatial intelligence is achieved with quantum speedup excludes that it is only the result of “ordinary” natural evolution (obeying the principle of causality). Indeed a quantum speedup in the development of spatial intelligence – and the consequent teleological character of this development – would imply the mutual causality in question.
(iii) The present notion of teleological evolution, namely of an evolution toward a goal with an attractor in the goal it will reach in the future, could also be seen in terms of a “systematically lucky” evolution, lucky from the standpoint of the objective of reaching its goal, indeed because attracted by it. Interestingly, also the notion of a systematically lucky evolution seems to have a precedent in pre-scientific thinking. It would be the Stoic notion of $\pi\rho o\nu o\iota\alpha$, foreknowledge $\left[44\right]$, sometimes also translated into providence and, in the present context, possibly translatable into advanced knowledge. Foreknowledge, or advanced knowledge, would have pushed the evolution toward its goal in a luckier and thus faster way than what is possible in the case of a causal evolution, exactly as the advanced knowledge about the solution does in quantum algorithms. This notion of $\pi\rho o\nu o\iota\alpha$ would also fit well with the notion of teleological evolution understood as a stochastic processes with transition probabilities time-symmetrically affected by the past and future of the process, with the luckier transitions more probable than in the classical case.
@@
(v) Michelangelo’s “Creation of Adam” on the ceiling of the Sistine Chapel could be the manifesto of the present theory. There is a mirror symmetry between creator and created $\left[45\right]$. Could this be interpreted as the time-symmetry – and the related mutual causation – between the fundamental physical laws with their physical constants since the beginning of time and the existence of sentient life today? Furthermore, the fact that the figures and shapes portrayed behind the creator figure would seem to be an anatomically accurate portrait of the human brain $\left[46\right]$, could it allude to the teleological character of the evolution of our intelligent species: the human brain being an attractor of the search for it? We would not feel like dismissing these analogies as sheer coincidences. As a matter of fact, in idealistic philosophy – in Schelling – aesthetic insight would reproduce the fundamental mechanisms of the production of life.
Overall, it would seem that the current ”scientific” theory of teleological evolutions reevaluates the idealistic side of philosophy against the materialistic side. However, the latter would also have its revenge, since a scientific justification of idealism could be ascribed to a materialistic view.
To sum up, it seems clear that the contrast between the causal and mutual causality views touches a deep mechanism in the history of thought. In an inevitably biased way, we would ascribe it to the unifying power of the current theory of teleological evolutions.
@@ 5 Research prospects
(ii) Look for toy models that capture the mechanism for which the mostly autonomous evolution of a subsystem, although decohered, inherits the quantum speedup and consequent teleological character of the evolution of a purely quantum system. Of course, here, the quantum system stands for the universe and the subsystem for the living Earth subsystem.
Note that the present notion for which the decohered behavior of a subsystem inherits quantum features from the behavior of the quantum system might apply to other cases. Possible examples are the life processes suspected to have higher than classical efficiency and consciousness suspected to need quantum nonlocality, the quantum system being here, of course, the universe.
@@
(II) Toy models of teleological Darwinian processes. These evolutions could be either approximated by pure quantum processes endowed with quantum speedup or described as decohered stochastic processes with transition probabilities time-symmetrically affected by the past and the future of the process.
@@ 6 Conclusions
We have trivially shown that the sheer existence of the quantum speedup of the optimal Grover quantum search algorithm logically implies the mutually exclusive $or$ of well defined causal loops. In each of them, it is as if the problem-solver knew in advance one of the possible halves of the information that specifies the solution of the problem she will produce and measure in the future and used this knowledge to produce the solution with correspondingly fewer computation steps. The number of computation steps needed to solve the problem in an optimal quantum way is that needed to solve it in an optimal logical way using the advanced knowledge in question.
On the one hand, being trivial, this claim is perfectly verifiable. On the other hand, a trivial demonstration of the existence of causal loops, a notion strongly questioned, should therefore be a given. This can be considered the fulcrum of the present work, and the source of its unconventionality. If, so simple, it has never been noticed outside the present retrocausal approach to the quantum speedup, it must be because a tacit no-retrocausality precept has prevented looking that way.
@@
Since quantum algorithms only involve the measurements of commuting observables, they are submitted to the same (classical) logic of the logical implication in question. Therefore, Grover algorithm must host the quantum superposition of the causal loops we are dealing with. However, of course, this cannot be in its quantum description that, being causal in character as that of any quantum process, cannot describe causal loops.
It should be noted that the usual quantum description of quantum algorithms presents another problem. It is limited to the process of solving the problem where the quantum speedup mathematically resides and is clearly incomplete. In fact, it consists of the unitary computation of the solution and the measurement of the latter. It lacks the initial measurement.
Remedying this incompleteness is another key point of the work. Extending the usual quantum description to the upstream process of setting the problem by measuring the problem-setting in the quantum superposition (or mixture) of all the possible problem-settings makes the quantum description of the quantum algorithm time-symmetric. The quantum description of Grover algorithm becomes that of a unitary evolution between an initial and a final one-to-one correlated measurement outcomes. Of course, being still causal, this extended quantum description cannot describe causal loops. However, time-symmetrizing it generates the quantum superposition of the causal loops logically implied by the sheer existence of the quantum speedup.
@@
We should require that the initial and final measurements time-symmetrically contribute to the selection of their pair of one-to-one correlated measurement outcomes. The selection performed by the initial measurement should propagate forward in time until becoming the state immediately before the final measurement. The selection performed by the final measurement should propagate backward in time until replacing the previous outcome of the initial measurement. This is a causal loops that changes its own past – the initial measurement outcome. It is indeed one of the causal loops logically implied by the sheer existence of the quantum speedup.
@@
Of course, we should take the quantum superposition of all the possible time-symmetrization instances, each for each possible way of time-symmetrically sharing the selection of the pair of measurement outcomes between the initial and final measurements. This leaves the quantum description of the unitary evolution between the initial and final measurement outcomes mathematically unaltered but changes the description of the structure of causality along that evolution. The single causal process of the ordinary quantum description is replaced by the quantum superposition of causal loops logically implied by the sheer existence of the quantum speedup. Being generated by a time-symmetrization, we call these causal loops time-symmetric.
@@
Note that the time-symmetrization in question turns out to be the only possible way of obtaining a quantum description of the quantum algorithm that describes the causal loops the quantum algorithm must host. It is therefore mandatory.
@@
We have shown that this immediately generalizes to any unitary evolution between two one-to-one correlated measurement outcomes that does not change the eigenvalue selected by the initial measurement (the problem-setting in quantum algorithms). The ordinary quantum description of any such process is incomplete and is completed by time-symmetrizing it. This leaves the ordinary quantum description of the unitary part of the quantum process mathematically unaltered but replaces its single causal process by a quantum superposition of time-symmetric causal loops.
@@
We believe that investigating this possibility is incumbent on the part of physics toward the natural sciences, after its having denied for centuries the possibility that teleological evolutions exist.
This immediately raises the problem of explaining the presence of quantum features (the quantum speedup) in evolutions of the living that can be the size of the Earth and have durations of billions of years. The obvious impossibility that the quantum features of the microscopic world scale up to such levels suggests the only possible solution: that they scale down from the most macroscopic and enduring evolution there is, that of the universe. Of course, at the same time we should assume, with quantum cosmology, that the evolution of the universe is unitary quantum. The fact that, once given the amount of resources involved, the quantum speedup is related to the time taken by the quantum process, under suitable conditions makes it inheritable by a subsystem (a natural evolution) of the quantum system (the universe).
@@
This is formally identical to a quantum search algorithm that unitarily builds up a quantum correlation between the value of the problem-setting since its beginning and the value of the solution and its end.
We have seen that any unitary evolution that builds up one-to-one correlated between an initial and a final measurement outcome without changing the eigenvalue selected by the initial measurement can occur with quantum speedup and does even if the search for that correlation is random. The latter is of course a conservative assumption; that there has been that building up by the evolution of the universe is what the Anthropic Principle tells us.
@@
In conclusion, we have seen that the ordinary, causal, quantum description of a class of processes involving quantum correlation is incomplete, exactly in the sense argued by Einstein and the others in their EPR paper, and is completed by time-symmetrizing it. This explains such diverse quantum phenomena as the quantum speedup and the consequent teleological character of quantum algorithms, quantum nonlocality, and, under a quantum cosmological interpretation of the Anthropic Principle, the teleological character of the evolutions of life. It also recovers a pre-scientific view of physical reality that had waned with the advent of modern science and its classical-causal view of physical phenomena.

### v7

@@ 3.1 Updating the EPR critique of the quantum description
Of course, the first assertion that the quantum description is incomplete goes back to the famous EPR paper by Einstein, Podolsky and Rosen $\left[29\right]$. As is well known, Einstein in particular, argued that the ordinary quantum description of quantum nonlocality is incomplete because it does not physically explain the “spooky action at a distance” that it mathematically predicts. They also stressed the need to devise ”hidden variables” that would complete the quantum description.
@@ 3.1.2 Generalization
In the first place, to be a quantum superposition of causal loop, the quantum algorithm does no more need to be optimal333 If the solution is not an invertible function of the problem setting, we can have one-to-one correlation between the initial and final measurement outcomes by requiring that the final measurement measures both the solution and the problem-setting. . The time-symmetrization of its ordinary quantum description always yields that it is as if the problem-solver knew in advance one of the possible halves of the information about the solution. In equivalent terms, there is a corresponding reduction of the computational complexity of the problem to be solved, and the dimension of the Hilbert space in which the search for the solution (or, identically, the one-to-one correlation) takes place. However, any of these facts just creates the room for a quantum speedup. To also have a quantum speedup, it takes that the problem-solver, or more generally the unitary evolution, exploits such a room. The room is fully exploited in the optimal quantum case, but it can also be not exploited at all or only partly exploited.
@@
By the way, as is well known, the EPR paper inspired works of great fundamental and practical interest like Bell’s inequality $\left[34\right]$ and Ekert’s quantum cryptography based on Bell’s theorem $\left[35\right]$. We do not understand why the argument concerning the incompleteness of the quantum description, central in that paper, has been removed by the scientific community. As is also well known, the question of the incompleteness of the quantum description has been at the center of the famous Einstein-Bohr dispute. Bohr, who was for the completeness, was eventually decreed the winner by the scientific community. Apparently that was once and for all.
@@ 3.2 Formalization
(2) This point is propaedeutic to point (4). It shows how to relativize with respect to the problem-solver the ordinary quantum description of the quantum algorithm, which is with respect to the customary observer external to the quantum process. Of course, the problem-solver must be concealed from the problem-setting selected by the initial measurement, which would give her full information about the solution of the problem444 We must assume that the problem-solver knows everything about the structure of the problem, thus also the relation between problem-setting and solution. .
@@ 3.2.1 Extending the usual quantum description of quantum algorithms to the process of setting the problem
We should start from the quantum state in the upper-left corner of Table I diagram and then follow the vertical and horizontal arrows. Initially, at time $t_{1}$, both registers, each by itself, are in a quantum superposition of all their basis vectors. By measuring $\hat{B}$ (the content of $B$) in this initial state, Bob selects the number of the drawer with the ball at random, say it comes out $01$ – see the state of register $B$ under the vertical arrow555 By the way, Bob would be free to unitarily change at will the sorted out number of the drawer with the ball. We omit this operation that would change nothing. .
@@
Note that this relativization relies on relational quantum mechanics $\left[36,37\right]$, where the quantum state is relative to the observer. In the following, we will continue to call the quantum description with respect to the external observer the ordinary quantum description of the quantum algorithm. Instead we will always specify that the other quantum description is that with respect to the problem-solver.
@@ 4 The possible physical basis of natural teleological evolutions
In the universe we are in, the search for the value of life becomes search for life, of course endowed with the same quantum speedup. We have also seen that, as a consequence, the evolution of life on Earth, obtainable from that of the universe by tracing out all the rest, even if decohered, should inherit the quantum speed up of the evolution of the universe toward life and would therefore be teleological too. The same would necessarily hold for the evolutions of species, etc..
@@ 4.2 Decohered teleological evolutions: what would they look like?
The above explanation is puzzling, of course. However, let us note that it is not more puzzling than the complete quantum description of the quantum computational speedup. We should keep in our mind that we are under the quantum cosmological assumption that should make macroscopic evolutions, the partial trace over all the rest of a quantum evolution of the universe, as puzzling as the microscopic quantum world.
@@ 4.3 How to model teleological natural evolutions
Let us note that, since the liquid water stage of it, the evolution of life on Earth should have been almost completely independent from the evolution of the rest of the universe. If it were completely independent, it would be factorizable out of the evolution of the universe, thus itself quantum and endowed with quantum speedup. Therefore, an obvious, approximate, way of modeling the decohered evolution of the living on Earth endowed with quantum speedup is just to ignore the fact it is decohered and model it in a quantum way. There are encouraging precedents to this. For example, it has been shown that modeling in a quantum way the interplay between species of predators competing for the same preys allows to explain the empirically ascertained species’ asymptotic behavior in a way otherwise difficult to explain $\left[38\right]$.
@@ 4.4 Impact on the natural sciences
This tension is captured by the following sentence by the evolutionary biologist Haldane $\left[39\right]$:
@@
No biologist, of course, does not mean no scientist. A paradigmatic example is the American philosopher Nagel $\left[40\right]$, who argues that the principles that explain the emergence of life should be teleological, rather than materialistic or mechanistic in character.
@@
Obviously, all this would go against the ban currently placed by the natural sciences on the teleological view.
Instead, it might go along with the fact that, in recent years, a group of evolutionary biologists $\left[41\right]$ expressed the need to rethink Darwin’s theory of evolution. That was after the discovery that some species’ evolution steps occurred far too quickly to be solely ascribed to random changes in the species’ DNA pool. Of course, the present teleological view of evolutions, with the associated quantum speedup, would offer a possible explanation.
It would also go along with the fact that some scientists have began to apply essentially quantum mechanical features like quantum entanglement and nonlocality to the macroscopic game of species, showing that this can explain some empirically ascertained features of it otherwise difficult to explain $\left[38\right]$.
@@ 4.5 Impact on philosophy
Of course, the current theory of teleological natural evolutions has a precise conceptual precedent in a very ancient notion of natural philosophy $\left[42\right]$. We should therefore be allowed to look for other possible cases where philosophical thinking has anticipated aspects of it.
@@
Let us call the present teleological view of evolutions and the causal view of today’s physics and the natural sciences respectively the mutual causality view666 Mutual causality between the initial and final measurement outcomes is what generates the quantum superposition of the time-symmetric causal loops. and the causal view.
The causal view and the mutual causality view seem to have respectively subtended two opposing school of thought throughout the history of philosophy. The causal view would underlie the materialist, mechanistic and reductionist views, while the mutual causality view would underlie the idealist, teleological and monist views respectively.
Consider the opposition between materialism and idealism. The former, about the observer being causally created by physical reality, is a causal view; the latter, about the mutual creation between observer and physical reality, like in the participatory Anthropic Principle of Wheeler $\left[32\right]$, and the anthropic cosmological principles of Barrow and Tipler $\left[12\right]$ is a mutual causality view.
@@
(i) For Plato, “Ideas [of our mind] are objective perfections that exist in themselves and for themselves and, at the same time, they are the causes of natural phenomena and constitute their unity” $\left[43\right]$. In modern language, the cause of natural phenomena that constitute their unity are the fundamental physical laws. The identity between the fundamental physical laws since the beginning of time and the ideas of our mind today is formally reminiscent of the identity (and mutual causality) between the number of the drawer with the ball (the problem-setting) randomly selected by Bob’s measurement at the beginning of the quantum algorithm and the number of the drawer with the ball (the solution of the problem) selected by Alice’s measurement at the end of it. It is as if the number of the drawer with the ball was determined by both the problem-setter and the problem-solver.
@@
The above should also go along with Fritjov Capra’s $\left[44\right]$ notion that there is identity between the fundamental states of consciousness described by eastern theosophies and the fundamental laws of quantum physics.
(ii) We do not need to be anthropocentric. This mutual causality would also be between fundamental physical laws and constants on the one side and spatial intelligence, often superior to our in non-human animals, on the other. The assumption that spatial intelligence is achieved with quantum speedup excludes that it is only the result of “ordinary” natural evolution (obeying the principle of causality). Indeed a quantum speedup in the development of spatial intelligence – and the consequent teleological character of this development – would imply the mutual causality in question.
(iii) The present notion of teleological evolution, namely of an evolution toward a goal with an attractor in the goal it will reach in the future, could also be seen in terms of a “systematically lucky” evolution, lucky from the standpoint of the objective of reaching its goal, indeed because attracted by it. Interestingly, also the notion of a systematically lucky evolution seems to have a precedent in pre-scientific thinking. It would be the Stoic notion of $\pi\rho o\nu o\iota\alpha$, foreknowledge $\left[45\right]$, sometimes also translated into providence and, in the present context, possibly translatable into advanced knowledge. Foreknowledge, or advanced knowledge, would have pushed the evolution toward its goal in a luckier and thus faster way than what is possible in the case of a causal evolution, exactly as the advanced knowledge about the solution does in quantum algorithms. This notion of $\pi\rho o\nu o\iota\alpha$ would also fit well with the notion of teleological evolution understood as a stochastic processes with transition probabilities time-symmetrically affected by the past and future of the process, with the luckier transitions more probable than in the classical case.
@@
(v) Michelangelo’s “Creation of Adam” on the ceiling of the Cappella Sistina could be the manifesto of the present theory. There is a mirror symmetry between creator and created $\left[46\right]$. Could this be interpreted as the time-symmetry – and the related mutual causation – between the fundamental physical laws with their physical constants since the beginning of time and the existence of sentient life today?
To sum up, it seems clear that the contrast between the causal and mutual causality views corresponds to an opposition that characterizes the entire history of thought. We would ascribe this unified view of the history of philosophy to the unifying power of the present “scientific” notion of teleological evolution that, under quantum cosmology, also extends to the macroscopic world.
@@ 5 Research prospects
(ii) Look for toy models that capture the mechanism for which the mostly autonomous evolution of a subsystem, although decohered, inherits the quantum speedup and the consequent teleological character of the evolution of a purely quantum system. Of course, here, the quantum system stands for the universe and the subsystem for the Earth subsystem.
Note that the present notion for which the decohered behavior of a subsystem inherits quantum features from the behavior of the quantum system might apply to other cases. Possible examples are the life processes suspected to have higher than classical efficiency and consciousness suspected to be endowed with quantum nonlocality, the quantum system being here, of course, a quantum universe.
@@
(II) Toy models of teleological Darwinian evolutions. These evolutions could be either approximated by pure quantum processes endowed with quantum speedup or described as decohered stochastic processes with transition probabilities time-symmetrically affected by the past and the future of the process.
Note that there might be a divide between the teleological view of evolution and the, of course dominant, causal/classical view. An evolution endowed with quantum speedup might not only mean a faster evolution; the mutual causality implicit in it could produce forms of evolution that are precluded to the causal/classical type.
@@ 6 Conclusions
We have (trivially) shown that the sheer existence of the quantum speedup of the optimal Grover quantum search algorithm logically implies the mutually exclusive $or$ of well defined causal loops. In each of them, it is as if the problem-solver knew in advance one of the possible halves of the information that specifies the solution of the problem she will produce and measure in the future and used this knowledge to produce the solution with correspondingly fewer computation steps. The number of computation steps needed to solve the problem in an optimal quantum way is that needed to solve it in an optimal logical way using the advanced knowledge in question.
@@
Since quantum algorithms only involve the measurements of commuting observables, they are submitted to the same (classical) logic of the logical implication in question. Therefore, Grover algorithm must host the quantum superposition of the causal loops we are dealing with – logically mutual exclusive possibilities go in quantum superpositions in the quantum world. However, of course, this cannot be in the quantum description of the quantum algorithm that, being causal, cannot describe causal loops.
It should be noted that we adopt a canonical quantum description of quantum processes, consisting of initial measurement, unitary evolution, and final measurement. In quantum algorithms, the initial measurement selects the problem-setting in a quantum superposition (or mixture) of all the possible problem-settings, the unitary evolution computes the corresponding solution, and the final measurement reads it.
Time-symmetrizing this, still causal, ordinary quantum description of the quantum algorithm generates the quantum superposition of the causal loops logically implied by the sheer existence of its quantum speedup.
@@
We require that the initial and final measurements time-symmetrically contribute to the selection of their pair of one-to-one correlated measurement outcomes; that the selection performed by the initial measurement propagates forward in time until becoming the state immediately before the final measurement and the selection performed by the final measurement propagates backward in time until replacing the previous outcome of the initial measurement. This is a causal loops that changes its own past – the initial measurement outcome. It is indeed one of the causal loops logically implied by the sheer existence of the quantum speedup.
@@
Of course, we should take the quantum superposition of all the possible time-symmetrization instances, each for each possible way of time-symmetrically sharing the selection of the pair of measurement outcomes between the initial and final measurements. This leaves the quantum description of the unitary evolution between the initial and final measurement outcomes mathematically unaltered but changes the behavior of causality along the evolution. The single causal process of the ordinary quantum description is replaced by the quantum superposition of causal loops logically implied by the sheer existence of the quantum speedup. Being generated by a time-symmetrization, we call these causal loops time-symmetric.
@@
Note that the time-symmetrization in question turns out to be the only possible way of obtaining a quantum description of the quantum algorithm with the quantum superposition of causal loops it must host. It is therefore mandatory.
@@
We have shown that this immediately generalizes to any unitary evolution between two one-to-one correlated measurement outcomes that does not change the eigenvalue selected by the initial measurement (the problem-setting in quantum algorithms). The ordinary quantum description of any such process is incomplete and is completed by time-symmetrizing it. Let us repeat ourselves: this leaves the ordinary quantum description of the unitary part of the quantum process mathematically unaltered but changes the description of the behavior of causality along that part. The single causal process of the ordinary quantum description is replaced by a quantum superposition of time-symmetric causal loops.
@@
It can be argued that investigating this possibility is incumbent on the part of the physical science toward the natural sciences, after its having asserted for centuries that teleological evolutions are unphysical.
This immediately raises the problem of explaining the presence of quantum features (the quantum speedup) in evolutions of the living that can be the size of the Earth and have durations of billions of years. The obvious impossibility that the quantum features of the microscopic world scale up to such levels suggests the only possible solution: that they scale down from the most macroscopic and enduring evolution there is, that of the universe. Of course, at the same time, we should assume, with quantum cosmology, that the evolution of the universe is unitary quantum.
@@
We have seen that any unitary evolution that builds up one-to-one correlation between an initial and a final measurement outcome without changing the eigenvalue selected by the initial measurement can occur with quantum speedup and does even if the search for that correlation is random. That the search for the correlation is random is of course a conservative assumption; that there has been this search by the evolution of the universe is what the Anthropic Principle says.
@@
In conclusion, we have seen that the ordinary, causal, quantum description of any unitary evolution between an initial and a final one-to-one correlated measurement outcomes that does not change the eigenvalue selected by the initial measurement is incomplete and is completed by time-symmetrizing it. This generates a quantum superposition of time-symmetric causal loops that explain such diverse quantum phenomena as the quantum speedup and the consequent teleological character of quantum algorithms, quantum nonlocality, and, under a quantum cosmological interpretation of the Anthropic Principle, the teleological character of the evolutions of life.
Although the time-symmetric causal loops we have been dealing with should be exempt from the paradoxes that plague the usual notion of causal loop insofar as they are imposed by logic and quantum mechanics, they remain puzzling for the consequences they entail. Being related to quantum correlation, they can also be thought to be pervasive. Once verified, their existence could introduce a paradigm shift in quantum mechanics.